Title

Sep 17 - 18

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: --

Summary

Stay closed due to bad weather conditions.

Sep 18-19

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: ---

Summary

At the sunset (?) fog, clouds and light rain. The telescope is at horizon for the leaking
in the dome, we can not perform calibration.

Sep 19-20

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: MODS

Summary

00:00 UT communication from Jeff Urban:
We will stay with the dome closed all the night for technical problem.
It to risky to unwrap the instruments from the cover, because there are chance of sudden rain and there
are not enough people during the night to cover them again.
Conditions at 00:00 UT: Humidity 97.9%, dew point 9.5\xB0C, temperature 10\xB0C

05:30 UT Conditions are better than expected. It is possible to open the dome

Data Summary

Project_RUN Objects Filter ExpTime Notes
FusiPecci L P04 dual grating 1 h COMPLETED
Standard Feige110 dual grating + blue grating 10 m  
FusiPecci M P09 dual grating 1 h COMPLETED
FusiPecci N P46 dual grating 45 m COMPLETED
FusiPecci P MCGC8 dual grating 10 m exposure stopped due to humidity
Calibrations        

Sep 20-21

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: MODS

Summary

At the beginning of the night the sky is really cloudy, humidity 93.8%, dewpoint 9.3\xB0C, temperature 7.9\xB0C, the TO is checking the conditions

Data Summary

Project_RUN Objects Filter ExpTime Notes
Calibrations        
FusiPecci O MCGC9 dual grating 1 h COMPLETED
FusiPecci P MCGC8 dual grating 30 m COMPLETED
Standard Feige110 dual grating 5 m  

Sep 21-22

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: ---

Summary

02.20 UT Due to a high umidity (96%) and a cloudy sky we stay closed

Sep 22-23

Observer: F. Cusano, S. Bisogni
Telescope Operator: S. Allanson
Support Astronomer: O. Khun
Instrument: ---

Summary

02:38 UT At the beginning of the night some clouds passing. Humidity 96%, dewpoint 9.1\xB0C, temperature 9.7\xB0C.
The TO opened the vent doors, but then closed them again due to humidity.

-- %USERSIG{FeliceCusano - 2014-11-10}%

Comments

 
Topic revision: r1 - 10 Nov 2014, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback